Two vectors of magnitude F and 2F have a resultant 1.5F. Calculate the angle between the vectors


1
Expert's answer
2020-07-22T10:35:15-0400

The Cosine Rule says

"(1.5F)^2=F^2+(2F)^2-2F(2F)\cos\theta"

Hence

"2.25=1+4-4\cos\theta"

"\cos\theta=0.6875"

"\theta=46.6^{\circ}"